a six
Relate probability to the study of genetics?
Probability is what chance something has to happen. The Punnett Square is a way how to predict in genetics how likely an offspring is to have a trait passed on from parents, or in other words find out the probability of a trait being in the phenotype or the genotype.
Well a Head Louse doesn't get very big about 3 Millimeters
20% (or 2 in 10 chance)
P(R on the third draw) = 3/5 = 0.60 = 60%
To calculate the probability of this event you have to consider all the possible outcomes that give a Red ball on the third draw. They are four:
[1R,2R,3R], [1R,2W,3R], [1W,2R,3R], [1W,2W,3R].
The probability for these four events are:
1.- P(1R,2R,3R) = (6/10)(5/9)(4/8) = 1/6
2.- P(1R,2W,3R) = (6/10)(4/9)(5/8) = 1/6
3.- P(1W,2R,3R) = (4/10)(6/9)(5/8) = 1/6
4.- P(1W,2W,3R) = (4/10)(3/9)(6/8) = 1/10
Then the probability looked for is the sum of all four:
P(R on the third draw) = 1/6 + 1/6 + 1/6 + 1/10 = 3/5 = 0.60 = 60%
How many face cards in a deck of 52 cards?
There are 4 jacks, 4 queens, and 4 kings or 12 face cards in a deck of 52 cards.
What are some advantages and disadvantages of multistage sampling?
a disadvantige is that a rocket gets disposed of after each flight
Give the example of why probabilities of A given B and B given A are not same?
Let's try this example (best conceived of as a squared 2x2 table with sums to the side). The comma here is an AND logical operator. P(A, B) = 0.1 P(A, non-B) = 0.4 P(non-A, B) = 0.3 P(non-A, non-B) = 0.2 then P(A) and P(B) are obtained by summing on the different sides of the table: P(A) = P(A, B) + P(A, non-B) = 0.1 + 0.4 = 0.5 P(B) = P(A,B) + P(non-A, B) = 0.1 + 0.3 = 0.4 so P(A given B) = P (A, B) / P (B) = 0.1 / 0.4 = 0.25 also written P(A|B) P(B given A) = P (A,B) / P (A) = 0.1 / 0.5 = 0.2 The difference comes from the different negated events added to form the whole P(A) and P(B). If P(A, non-B) = P (B, non-A) then P(A) = P(B) and also P(A|B) = P(B|A).
This is a sampling problem where the entire population is selected and there is no replacement. The chance of selecting all cities in alphabetical order is: (1/10)*(1/9)*(1/8)*(1/7)*(1/6)*(1/5)*(1/4)*(1/3)*(1/2) of 1/(10!) = 2.75 e-7. If we consider alphabetical order can be highest to lowest or lowest to highest, then the probability doubles or: 2/(10!)
What are the characteristics of a normal distribution curve?
Characteristics of a Normal Distribution
1) Continuous Random Variable.
2) Mound or Bell-shaped curve.
3) The normal curve extends indefinitely in both directions, approaching, but never touching, the horizontal axis as it does so.
4) Unimodal
5) Mean = Median = Mode
6) Symmetrical with respect to the mean
That is, 50% of the area (data) under the curve lies to the left of
the mean and 50% of the area (data) under the curve lies
to the right of the mean.
7) (a) 68% of the area (data) under the curve is within one
standard deviation of the mean
(b) 95% of the area (data) under the curve is within two
standard deviations of the mean
(c) 99.7% of the area (data) under the curve is within three
standard deviations of the mean
8) The total area under the normal curve is equal to 1.
Is a parallelogram's diagonals perpendicular?
If the parallelogram happens to also be a rhombus (i.e. has all sides equal in length) then yes, otherwise no.
Definition OC Curves or Operating Characteristic Curves refer to a graph of attributes of a sampling plan considered during management of a project which depicts the percent of lots or batches which are expected to be acceptable under the specified sampling plan and for a specified process quality. The specified sampling plan may be singular, sequential or iterative and may be using a particular size of a sample depending upon the demands of the project and could yield the results of acceptance or rejection based on a specified criteria. It helps in the selection of sampling plans It aids in the selection of plans that are effective in reducing risks. It can help in keeping the high cost of inspection low. There are two ways of calculating the OC curves. Binomial Distribution Poisson Formula Type A - Gives the probability of acceptance for an individual lot coming from finite production. Type B - Gives the probability of acceptance for lots coming from a continuous process. Type C - Gives the long-run percentage of product accepted during the sampling phase.
How is Probability used in the gambling industry?
It depends on what you mean, but here is a basic description of how it works. Payouts in gambling games are based on the probability of winning, but the payout is less than the true probability of winning. Just as a basic example, the odds of rolling a 12 in the game Craps is 1 in 36, but if bet on the 12 and win, you only win 30 times the amount bet. So basically, you would have to make a $1 bet 36 times to win $30.
How are quadratic equations used in real life?
Many real life physics problems are parabolic in nature. Parabolas can be shown as a quadratic equation. If you have two variables then usually you can use the equation to find the best solution to a problem.
Also, it is a beginning in the world of mathematical optimization. Some equations use more than two variables and require the technique used to solve quadratics to solve them. I just ran an optimization of 128 variables. To understand the parameters I needed to set I had to understand quadratics.
Since we want both a head and a 6 we need the multiplication rule of probabilities. Chance of head is 1/2 and chance of 6 is 1/6 so chance of both is 1/12. P( head and 6)= 1/12= about .08333
What is the probability of getting five tails in a row?
The probability of getting five tails in a row is 1/2^5, or 1 in 32.
The probability of getting five heads in a row is 1/2^5, or 1 in 32.
Thus, the probability of getting either five heads or five tails in five tosses is 1 in 16.
(The caret symbol means "to the power of," as in 2^5 means "2 to the 5th power.")
If one card is drawn from a regular desk cards what is the probability the card will not be a ace?
There are total of 52 cards in a deck and 4 of those are Aces. Therefore, there are 48 card non-ace. The probability of drawing one card that is not an ace will be 48 divided by 52. The answer is 12/13.
What does the Probability of success in a binomial trial symbol?
In typical notation, "p" is the probability of sucess and "q" is the probability of failure. So q = 1 - p. But for your question: p = p.
15/100=0.15
http://wiki.answers.com/Q/The_letter_of_the_word_statistics_are_scrambled_and_the_letters_are_rearranged_find_the_probability_that_the_vowels_occur_together"
P(0 < Z < z) = 0.3770 is the same as looking at P(Z < z) = 0.8770 because the other half of the curve (anything less than 0) has probability of 0.5. Now this is a problem of just looking it up from the table. The table gives a value z = 1.16 for the probability of 0.8770. So P(0 < Z < 1.16) = 0.3770.
What is the probability of drawing a Queen of Clubs or a Queen of Spades from a deck of 52 cards?
Reason:
There are only 2 cards in the deck of 52 which could be either a Queen of Clubs or Queen or Spades, So there is a 2 out of 52 chance, or 3.8461538 % chance.